OpenStudy (anonymous):

The area of a right triangle with one leg measuring 5 ft. and hypotenuse measuring 13 ft. is

OpenStudy (anonymous):



Parth (parthkohli):

We know that 5,12,13 are the sides according to the Pythagorean. \(\Large \color{Black}{\Rightarrow 0.5bh = Area(Triangle) }\) Base = 5 Height = 12

OpenStudy (anonymous):

Knowing the hypotenuse is 13 ft and one side of the triangle is 5 ft. You can relate them together by Pythagoras Theorem. 13^2 = 5^2 + x^2 , where x is the length of the unknown side. x^2 = 144 x= 12 Given the length of the 2 sides that are perpendicular to each other, you can easily find the area by (area = 0.5 * base * height) area = 0.5 * 12 * 5 =30 ft ^2

OpenStudy (anonymous):

thx

OpenStudy (feemeegaming):

teehee